What is the percentage error in the reading of a voltmeter that shows 239 volts for an


actual 245 volts?

1
Expert's answer
2021-11-13T01:28:40-0500

Absolute error

"=|Actual- Measured \>value|\\=|239-245|=6"

"Percentage \>error=\frac{Absolute\> error}{Actual\>value}×100\%"

"=\frac{6}{245}×100\%=2.45\%"